OpenStudy (briensmarandache):

solve 3sine x -4sine^3 x = sqrt(3)/2

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ 3\sin(x) -4\sin^3(x)=\frac{\sqrt{3}}{2} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \sin(x)(3 -4\sin^2x)=\frac{\sqrt{3}}{2} }\) can you solve the first part when it is equivalent to sin(x)?

OpenStudy (briensmarandache):

i got that far but im not sure how to continue

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \sin(x)(3 -4\sin^2x)=\frac{\sqrt{3}}{2} }\) \(\Large\color{black}{ \sin(x)=\frac{\sqrt{3}}{2} }\) or \(\Large\color{black}{ 3 -4\sin^2x=\frac{\sqrt{3}}{2} }\)

OpenStudy (solomonzelman):

can you solve the first, the left part?

OpenStudy (briensmarandache):

for some reason i was to further factor 3-4sine^2x

OpenStudy (briensmarandache):

so x = arcsine(sqrt3/2), arcsine (+/- sqrt( ((sqrt3/2) -3) / 4)

OpenStudy (solomonzelman):

you can't find the x, sine of which is sqrt{3}/2, is that your trouble? Draw a 30-60-90 \(\large\color{black}{ \triangle }\).

OpenStudy (briensmarandache):

beacuse the hyp is smaller than opp

OpenStudy (solomonzelman):